OpenStudy (anonymous):

An auto manufacturer estimates that its lowest sport utility vehicle will travel 336 highway miles on one tank of gas. If the sport utility vehicles gas tank holds 24 gallons, how far can a driver expect to travel on 16 gallons?

OpenStudy (texaschic101):

Set it up as a proportion... 24/336 = 16/x (24 gallons goes 336 miles = 16 gallons goes x miles) now cross multiply... (24)(x) = (336)(16) 24x = 5376 x = 5376/24 x = 224 16 gallons will take you 224 miles. Do you have any questions ?